# Hyperbolic regularization effects for degenerate elliptic equations

### v1

@@ 6.2. Step 2: Reduction to degenerate curves
Thanks to the unique continuation result of [19], see [37, Proposition 3.1] and Remark 6.2, $Dw$ is either constant in $\Omega$ or takes values into the degenerate part $\Gamma_{d}=\gamma(\{\det(\gamma^{\prime})=0\})$. We can assume that we are in the latter case. According to Theorem 1.3, $Dw$ takes values at almost every point in one single connected component $\mathcal{C}$ of $\Gamma_{d}$. The required interval is then $I_{*}\doteq\gamma^{-1}(\mathcal{C})$.∎
@@
Note that [37, Proposition 3.1] is stated for $C^{2}$ curves. The results of [37] are restricted to $C^{2}$ curves for other reasons, but [37, Proposition 3.1] relies on [19], which does not require any smoothness, and on [36], where the proof is written for smooth curves, but works for $C^{1}$ curves modulo minor adaptations.
@@ 6.3. Step 3: Degenerate graphical curves
for all $v\in\mathbb{R}$. This is the kinetic formulation [39] associated with the scalar conservation law (6.8), in the case of zero entropy production. The fact that $f$ cannot be affine on any open interval ensures that for any $v_{1}<v_{2}\in[a,b]$ one can find $v_{1}<\tilde{v}_{1}<\tilde{v}_{2}<v_{2}$ such that $(1,f^{\prime}(\tilde{v}_{1}))$ and $(1,f^{\prime}(\tilde{v}_{2}))$ are linearly independent in $\mathbb{R}^{2}$. This property, combined with the free transport equation, implies that $u$ is continuous, see e.g. [17, Proposition 6]. We sketch here the argument for the readers’ convenience: to prove e.g. upper semicontinuity it suffices to show, for all $v_{1}<v_{2}$, the existence of $0<\delta<1$ such that
@@
Let $L=\|\partial_{x}h\|_{L^{\infty}(\Omega)}$ and $C=\max_{[-L,L]}|f^{\prime}|\,.$ To show that $u$ is an entropy solution it suffices to do so in a neighborhood of any $(t_{0},x_{0})\in\Omega$. Pick $T>0$ such that $t_{0}-T/2,t_{0}+T/2]\times[x_{0}-CT,x_{0}+CT]\subset\Omega$. Translating the coordinates, we assume $t_{0}-T/2=0$ and $x_{0}=0$. We choose a bounded, compactly supported, $L$-Lipschitz function $\tilde{h}_{0}\colon\mathbb{R}\to\mathbb{R}$ such that $\tilde{h}_{0}(x)=h(0,x)$ for $x\in[-CT,CT]$, and we consider the unique viscosity solution $\tilde{h}\colon[0,T]\times\mathbb{R}\to\mathbb{R}$ of (6.11) in $0,T]\times\mathbb{R}$, see for instance [10, Theorem VI.2]. Thanks to the local comparison principle [10, Theorem V.3] it coincides with $h$ in the cone $\mathcal{C}=\{0\leq t\leq T,\,|x|\leq C(T-t)\}\,.$ Moreover $\tilde{h}$ can be obtained as $\tilde{h}=\lim\tilde{h}_{\varepsilon}$, in the sense of distributions, of the vanishing viscosity approximation $\tilde{h}_{\varepsilon}$ solving
@@
with $\tilde{u}_{0}=\partial_{x}\tilde{h}_{0}$. Thanks to [12, § 6.3], we have $\tilde{u}_{\varepsilon}\to\tilde{u}$ as distributions, where $\tilde{u}$ is the unique entropy solution such that $\tilde{u}(0,\cdot)=\tilde{u}_{0}$. Moreover we also have $\tilde{u}_{\varepsilon}\to\partial_{x}\tilde{h}$ as distributions, and since $\tilde{h}=h$ in $\mathcal{C}$ we deduce that $\tilde{u}=u$ in $\mathcal{C}$, so $u$ is an entropy solution in $\mathcal{C}$, and therefore in a neighborhood of $(t_{0},x_{0})$. ∎
@@ 6.4. Step 4: Partial regularity
The continuity of $\Psi$ is contained in [37, Lemma 5.1], we provide the proof for completeness. Without loss of generality we assume that $\gamma$ is an arc-length parametrization, that is, $|\gamma^{\prime}|=1$ on $I$. This, together with the degeneracy $\det(\gamma^{\prime})=0$, implies that $\mathfrak{a}=[\gamma^{\prime}]_{\mathcal{H}}$ and $\mathfrak{b}=[\gamma^{\prime}]_{\bar{\mathcal{H}}}$ satisfy $|\mathfrak{a}|=|\mathfrak{b}|=1/2$, compare (2.3). As continuous maps from $I$ to $\frac{1}{2}\mathbb{S}^{1}$, they can be written as $2\mathfrak{a}=e^{i\alpha}$, $2\mathfrak{b}=e^{i\beta}$, for some liftings $\alpha,\beta\colon I\to\mathbb{R}$. These liftings are continuous in the case $I=[a,b]$, but may fail to be periodic in the non-simply-connected case $I=\mathbb{R}/L{\mathbb{Z}}$, so that they may not be continuous maps from $I$ to $\mathbb{R}$. However, in this case they can still be identified with continuous functions on $\mathbb{R}$ which satisfy $\alpha(t+L)=\alpha(t)+2k\pi$, $\beta(t+L)=\beta(t)+2\ell\pi$ for some $k,\ell\in{\mathbb{Z}}$. Direct calculation then shows that
@@
Thanks to Theorem A and Theorem B, it suffices to show that the set $\operatorname{Reg}(w)$ defined in Definition 1.2 is open. Let $\varepsilon>0$ be as in Proposition 6.4. Thanks to Theorem A, if $x\in\operatorname{Reg}(w)$ then there exists $\delta=\delta(\varepsilon)>0$ such that $\operatorname{diam}([Dw](B_{\delta}(x)))\leq\varepsilon$, so $w$ is $C^{1}$ in $B_{\delta}(x)$, hence $B_{\delta}(x)\subset\operatorname{Reg}(w)$. ∎
@@ 6.5. Step 5: Entropy productions
In this section and the next one, the goal is to analyze the structure of the solution $w$ to $Dw\in\Gamma$ a.e. around singular points, thus completing the proof of Theorem D. Recall that we work under assumption (6.4) for $\Gamma$.
Heuristically, one can expect a rigid structure of $w$ near singularities because of the local constancy of $Dw$ along characteristic lines outside $\operatorname{Sing}(w)$ combined with the fact that generic lines do not intersect the $\mathcal{H}^{1}$-negligible singular set $\operatorname{Sing}(w)$, i.e. Propositions 6.4-6.4. One difficulty is that characteristic lines are not generic since their direction depends on the value of $Dw$, so it is not at all obvious that their typical behavior is not to intersect $\operatorname{Sing}(w)$. This difficulty can be overcome by taking inspiration from the kinetic formulation (6.10) of the scalar conservation law (6.8) used in the proof of Proposition 6.3. There, the kinetic variable $v$ is decoupled from the value of the solution $u(x)$, and for generic $v$ the line $x+\mathbb{R}(1,f^{\prime}(v))$ will not intersect a $\mathcal{H}^{1}$-negligible set. Since the kinetic formulation (6.10) is related to the vanishing of entropy productions it becomes natural to introduce similar tools here. We refer the reader to [50] for a systematic treatment of the link between entropies and kinetic formulations of conservation laws.
@@
Applying this to $A=\Sigma_{1}\circ\gamma$ and $B=\Sigma_{2}\circ\gamma$, and noting that $\Sigma(Dw)=\Sigma(\gamma(u))$, we get $\operatorname{div}\Sigma(Dw)=0$ in $B_{r}(x)$. This is valid for any $x\in\Omega\setminus\operatorname{Sing}(w)$ and some $r=r(x)>0$, hence $\operatorname{div}\Sigma(Dw)=0$ in $\mathcal{D}^{\prime}(\Omega\setminus\operatorname{Sing}(w))$. Since $\Sigma(Dw)\in L^{\infty}(\Omega)$ and $\mathcal{H}^{1}(\operatorname{Sing}(w))=0$, [27, Theorem 4.1(b)] yields $\operatorname{div}\Sigma(Dw)=0$ in $\mathcal{D}^{\prime}(\Omega)$. ∎
Through this Proposition, we could in principle apply the strategy of [37, §7] to conclude the proof. However, in [37] the curve $\Gamma$ is assumed to have $C^{2}$ regularity, and some nontrivial adaptations are required to deal with our lower $C^{1}$ regularity.
@@ 6.5.1. The case of a nonclosed curve
Recall from Lemma 6.4 that $\Psi$ is not constant on any open interval. Thus, for any $\alpha_{1}<\alpha_{2}\in[a,b]$, there exist $\alpha_{1}<\tilde{\alpha}_{1}<\tilde{\alpha}_{2}<\alpha_{2}$ such that $\Psi(\tilde{\alpha}_{1})$ and $\Psi(\tilde{\alpha}_{2})$ are linearly independent in $\mathbb{R}^{2}$. Combining this with (6.14), we conclude that $\theta$ is continuous by [17, Proposition 6] (whose argument is recalled in the proof of Proposition 6.3), and hence so is $Dw=\gamma(\theta)$. ∎
@@ 6.5.2. The case of a closed curve
Lemma 6.5.2 allows us to use [25, Theorem 6.5], which we recall for the reader’s convenience.
@@
Here and in what follows we implicitly identify a map $v\colon\Omega\to\mathbb{RP}^{1}=\mathbb{S}^{1}/\{\pm 1\}$ with any lifting $\tilde{v}\colon\Omega\to\mathbb{S}^{1}$ such that $v=\{\pm\tilde{v}\}$. In [25, Theorem 6.5], this result is stated for the map $\tilde{v}^{2}$ which uniquely determines $v$ and still satisfies (6.15). Theorem 6.15 is a direct translation of that statement, taking into account the aforementioned implicit identification.
@@ 7. Solutions of degenerate equations
Let $G\colon\mathbb{R}^{2}\to\mathbb{R}^{2}$ a continuous and strictly monotone vector field and let $u\in\operatorname{Lip}(B_{1})$ be a solution of (1.1). In [34] it is shown that, for $x_{0}\in B_{1}$, any blow-up limit $u_{\infty}\in\mathcal{B}(u)(x_{0})$ is either affine or satisfies
$Du_{\infty}\in\mathcal{D}=\mathcal{D}_{-}\cap\mathcal{D}_{+}\quad\text{a.e.,}$
where $\mathcal{D}_{\pm}$ are defined in (1.16). Recalling from Proposition 2.2 the correspondence between (1.1) and (1.8), there is a direct reformulation in terms of differential inclusions.
@@
Let $K\subset\mathbb{R}^{2\times 2}$ satisfy (1.9), $w\in\operatorname{Lip}(B_{1},\mathbb{R}^{2})$ solve $Dw\in K$ a.e. in $B_{1}$, and $x_{0}\in B_{1}$. Any blow-up limit $w_{\infty}\in\mathcal{B}(w)(x_{0})$ is either affine or satisfies $Dw_{\infty}\in\mathcal{K}_{*}$ a.e., where
@@
In this section, we combine this property of blow-up limits with Theorems A-B-C-D and structural assumptions on the degenerate set $\mathcal{D}$, or equivalently $\mathcal{K}_{*}$, in order to deduce partial regularity properties of (1.8), or equivalently (1.1), obtaining in particular Theorem E.
@@ 7.1. Regularity threshold
Thanks to Theorem A and Theorem B, it suffices to show that the set $\operatorname{Reg}(w)$ defined in Definition 1.2 is open. Let $\varepsilon=\varepsilon_{*}(K)/2>0$. By Theorem A, if $x\in\operatorname{Reg}(w)$ then there exists $\delta>0$ such that $\operatorname{diam}([Dw](B_{\delta}(x)))\leq\varepsilon$, so $w$ is $C^{1}$ in $B_{\delta}(x)$ by definition (7.2) of $\varepsilon_{*}(K)$, hence $B_{\delta}(x)\subset\operatorname{Reg}(w)$. ∎
Thanks to Theorem 7.1 and Theorem A, we can relate the regularity threshold $\varepsilon_{*}$ of $K$ with the regularity thresholds of the connected components of $\mathcal{K}_{*}$.
@@
If the infimum is zero, there is nothing to show. Assume therefore that it is positive, and fix $0<\varepsilon<\inf\{\varepsilon_{*}(\mathcal{C})\}$. Let $w\colon B_{1}\to\mathbb{R}^{2}$ satisfy $Dw\in K$ a.e. and $\operatorname{diam}([Dw](B_{1}))\leq\varepsilon$. This pointwise constraint is preserved under blow-up: for any $x_{0}\in B_{1}$ and $w_{\infty}\in\mathcal{B}(w)(x_{0})$ we have $\operatorname{diam}([Dw_{\infty}])(B_{1})\leq\varepsilon$. By Theorem 7.1, the blow-up limit $w_{\infty}$ is either affine or satisfies $Dw_{\infty}\in\mathcal{K}_{*}$ a.e., hence $Dw_{\infty}\in\mathcal{C}$ a.e. for some connected component $\mathcal{C}$ of $\mathcal{K}_{*}$ by Theorem 1.3. By definition of $\varepsilon_{*}(\mathcal{C})$ this implies $w_{\infty}\in C^{1}(B_{1},\mathbb{R}^{2})$, hence $x_{0}\in\operatorname{Reg}(w)$ by Definition 1.2. Thus $\operatorname{Reg}(w)=B_{1}$ and $\varepsilon_{*}(K)\geq\varepsilon$. ∎
@@
Under this assumption, every connected component $\mathcal{C}$ of $\mathcal{K}_{*}$ satisfies $\mathcal{C}\subset\Gamma=\gamma(I)$ for some $I=[a,b]$ or $\mathbb{R}/L{\mathbb{Z}}$ and $\gamma\colon I\to\Gamma$ a $C^{1}$ homeomorphism with $|\gamma^{\prime}|>0$. Since $\mathcal{C}$ is compact and connected and $\gamma$ is a homeomorphism, we must in fact have $\mathcal{C}=\gamma(J)$ for some compact interval $J\subset I$. So assumption (7.3) is actually equivalent to every connected component $\mathcal{C}$ of $\mathcal{K}_{*}$ being a compact $C^{1}$ curve, which implies $\varepsilon_{*}(\mathcal{C})>0$ by Proposition 6.4. In fact, we will see in the next §7.2 that under assumption (7.3), the property $\varepsilon_{*}(K)>0$ implies a much stronger conclusion than that of Lemma 7.1.
@@
This can only happen if $a=b$, and hence we conclude that $h$ fulfills (2.10). Since $h$ is one-homogeneous, we have that $w_{\alpha}(z)\doteq\alpha w(z)$ solves (2.8) for $h$ as in (7.4) for all $\alpha\in\mathbb{R}$. Moreover $\operatorname{Sing}(w_{\alpha})=\{0\}$, regardless of how small $\alpha$ is, showing that no $\varepsilon$-regularity theorem can hold for the PDE (2.8) for such $h$. In other words, the regularity threshold (7.2) of the set $K=\{(a,h(a))\colon|a|\leq 1\}$ is $\varepsilon_{*}(K)=0$. The problem in this case is, in fact, fully degenerate: for $h$ as in (7.4), passing to the equivalent formulation in terms of a monotone field $G$ as in (1.1) (see Propositions 2.2-2.2), we find that $\mathcal{D}_{+}\cap\mathcal{D}_{-}=\mathbb{\mathbb{R}}^{2}$, or equivalently $K=\mathcal{K}_{*}$.
7.2. Characterization of blow-up limits
Since blow-up limits $w_{\infty}$ are entire solutions of $Dw_{\infty}\in\mathcal{C}$ a.e. for some connected component $\mathcal{C}$ of $\mathcal{K}_{*}$, under assumption (7.3) we know by Theorem D that they are $C^{1}$ away from at most two singular points. This statement is valid for any entire solution, but blow-up limits have the additional property that their Hessian determinant vanishes away from the origin by Theorem C. Using this, we establish that they have in fact at most one singularity.

### v2

@@ 8.2. Step 2: Reduction to degenerate curves
Thanks to the unique continuation result of [21], see [39, Proposition 3.1] and Remark 8.2, $Dw$ is either constant in $\Omega$ or takes values into the degenerate part $\Gamma_{d}=\gamma(\{\det(\gamma^{\prime})=0\})$. We can assume that we are in the latter case. According to Theorem 2.4, $Dw$ takes values at almost every point in one single connected component $\mathcal{C}$ of $\Gamma_{d}$. The required interval is then $I_{*}\doteq\gamma^{-1}(\mathcal{C})$.∎
@@
Note that [39, Proposition 3.1] is stated for $C^{2}$ curves. The results of [39] are restricted to $C^{2}$ curves for other reasons, but [39, Proposition 3.1] relies on [21], which does not require any smoothness, and on [38], where the proof is written for smooth curves, but works for $C^{1}$ curves modulo minor adaptations.
@@ 8.3. Step 3: Degenerate graphical curves
for all $v\in\mathbb{R}$. This is the kinetic formulation [41] associated with the scalar conservation law (8.8), in the case of zero entropy production. The fact that $f$ cannot be affine on any open interval ensures that for any $v_{1}<v_{2}\in[a,b]$ one can find $v_{1}<\tilde{v}_{1}<\tilde{v}_{2}<v_{2}$ such that $(1,f^{\prime}(\tilde{v}_{1}))$ and $(1,f^{\prime}(\tilde{v}_{2}))$ are linearly independent in $\mathbb{R}^{2}$. This property, combined with the free transport equation, implies that $u$ is continuous, see e.g. [19, Proposition 6]. We sketch here the argument for the readers’ convenience: to prove e.g. upper semicontinuity it suffices to show, for all $v_{1}<v_{2}$, the existence of $0<\delta<1$ such that
@@
Let $L=\|\partial_{x}h\|_{L^{\infty}(\Omega)}$ and $C=\max_{[-L,L]}|f^{\prime}|\,.$ To show that $u$ is an entropy solution it suffices to do so in a neighborhood of any $(t_{0},x_{0})\in\Omega$. Pick $T>0$ such that $t_{0}-T/2,t_{0}+T/2]\times[x_{0}-CT,x_{0}+CT]\subset\Omega$. Translating the coordinates, we assume $t_{0}-T/2=0$ and $x_{0}=0$. We choose a bounded, compactly supported, $L$-Lipschitz function $\tilde{h}_{0}\colon\mathbb{R}\to\mathbb{R}$ such that $\tilde{h}_{0}(x)=h(0,x)$ for $x\in[-CT,CT]$, and we consider the unique viscosity solution $\tilde{h}\colon[0,T]\times\mathbb{R}\to\mathbb{R}$ of (8.11) in $0,T]\times\mathbb{R}$, see for instance [12, Theorem VI.2]. Thanks to the local comparison principle [12, Theorem V.3] it coincides with $h$ in the cone $\mathcal{C}=\{0\leq t\leq T,\,|x|\leq C(T-t)\}\,.$ Moreover $\tilde{h}$ can be obtained as $\tilde{h}=\lim\tilde{h}_{\varepsilon}$, in the sense of distributions, of the vanishing viscosity approximation $\tilde{h}_{\varepsilon}$ solving
@@
with $\tilde{u}_{0}=\partial_{x}\tilde{h}_{0}$. Thanks to [14, § 6.3], we have $\tilde{u}_{\varepsilon}\to\tilde{u}$ as distributions, where $\tilde{u}$ is the unique entropy solution such that $\tilde{u}(0,\cdot)=\tilde{u}_{0}$. Moreover we also have $\tilde{u}_{\varepsilon}\to\partial_{x}\tilde{h}$ as distributions, and since $\tilde{h}=h$ in $\mathcal{C}$ we deduce that $\tilde{u}=u$ in $\mathcal{C}$, so $u$ is an entropy solution in $\mathcal{C}$, and therefore in a neighborhood of $(t_{0},x_{0})$. ∎
@@ 8.4. Step 4: Partial regularity
The continuity of $\Psi$ is contained in [39, Lemma 5.1], we provide the proof for completeness. Without loss of generality we assume that $\gamma$ is an arc-length parametrization, that is, $|\gamma^{\prime}|=1$ on $I$. This, together with the degeneracy $\det(\gamma^{\prime})=0$, implies that $\mathfrak{a}=[\gamma^{\prime}]_{\mathcal{H}}$ and $\mathfrak{b}=[\gamma^{\prime}]_{\bar{\mathcal{H}}}$ satisfy $|\mathfrak{a}|=|\mathfrak{b}|=1/2$, compare (3.3). As continuous maps from $I$ to $\frac{1}{2}\mathbb{S}^{1}$, they can be written as $2\mathfrak{a}=e^{i\alpha}$, $2\mathfrak{b}=e^{i\beta}$, for some liftings $\alpha,\beta\colon I\to\mathbb{R}$. These liftings are continuous in the case $I=[a,b]$, but may fail to be periodic in the non-simply-connected case $I=\mathbb{R}/L{\mathbb{Z}}$, so that they may not be continuous maps from $I$ to $\mathbb{R}$. However, in this case they can still be identified with continuous functions on $\mathbb{R}$ which satisfy $\alpha(t+L)=\alpha(t)+2k\pi$, $\beta(t+L)=\beta(t)+2\ell\pi$ for some $k,\ell\in{\mathbb{Z}}$. Direct calculation then shows that
@@
Thanks to Theorem 6.2, it suffices to show that the set $\operatorname{Reg}(w)$ defined in Definition 6 is open. Let $\varepsilon>0$ be as in Proposition 8.4. Thanks to Theorem 4.1, if $x\in\operatorname{Reg}(w)$ then there exists $\delta=\delta(\varepsilon)>0$ such that $\operatorname{diam}([Dw](B_{\delta}(x)))\leq\varepsilon$, see the proof of Theorem 6.2 for a similar argument, so $w$ is $C^{1}$ in $B_{\delta}(x)$, hence $B_{\delta}(x)\subset\operatorname{Reg}(w)$. ∎
@@ 8.5. Step 5: Entropy productions
In this section and the next one, the goal is to analyze the structure of the solution $w$ to $Dw\in\Gamma$ a.e. around singular points, thus completing the proof of Theorem 8.1. Recall that we work under assumption (8.4) for $\Gamma$.
Heuristically, one can expect a rigid structure of $w$ near singularities because of the local constancy of $Dw$ along characteristic lines outside $\operatorname{Sing}(w)$ combined with the fact that generic lines do not intersect the $\mathcal{H}^{1}$-negligible singular set $\operatorname{Sing}(w)$, i.e. Propositions 8.4-8.4. One difficulty is that characteristic lines are not generic since their direction depends on the value of $Dw$, so it is not at all obvious that their typical behavior is not to intersect $\operatorname{Sing}(w)$. This difficulty can be overcome by taking inspiration from the kinetic formulation (8.10) of the scalar conservation law (8.8) used in the proof of Proposition 8.3. There, the kinetic variable $v$ is decoupled from the value of the solution $u(x)$, and for generic $v$ the line $x+\mathbb{R}(1,f^{\prime}(v))$ will not intersect a $\mathcal{H}^{1}$-negligible set. Since the kinetic formulation (8.10) is related to the vanishing of entropy productions it becomes natural to introduce similar tools here. We refer the reader to [52] for a systematic treatment of the link between entropies and kinetic formulations of conservation laws.
@@
Applying this to $A=\Sigma_{1}\circ\gamma$ and $B=\Sigma_{2}\circ\gamma$, and noting that $\Sigma(Dw)=\Sigma(\gamma(u))$, we get $\operatorname{div}\Sigma(Dw)=0$ in $B_{r}(x)$. This is valid for any $x\in\Omega\setminus\operatorname{Sing}(w)$ and some $r=r(x)>0$, hence $\operatorname{div}\Sigma(Dw)=0$ in $\mathcal{D}^{\prime}(\Omega\setminus\operatorname{Sing}(w))$. Since $\Sigma(Dw)\in L^{\infty}(\Omega)$ and $\mathcal{H}^{1}(\operatorname{Sing}(w))=0$, [29, Theorem 4.1(b)] yields $\operatorname{div}\Sigma(Dw)=0$ in $\mathcal{D}^{\prime}(\Omega)$. ∎
Through this Proposition, we could in principle apply the strategy of [39, §7] to conclude the proof. However, in [39] the curve $\Gamma$ is assumed to have $C^{2}$ regularity, and some nontrivial adaptations are required to deal with our lower $C^{1}$ regularity.
@@ 8.5.1. The case of a nonclosed curve
Recall from Lemma 8.4 that $\Psi$ is not constant on any open interval. Thus, for any $\alpha_{1}<\alpha_{2}\in[a,b]$, there exist $\alpha_{1}<\tilde{\alpha}_{1}<\tilde{\alpha}_{2}<\alpha_{2}$ such that $\Psi(\tilde{\alpha}_{1})$ and $\Psi(\tilde{\alpha}_{2})$ are linearly independent in $\mathbb{R}^{2}$. Combining this with (8.14), we conclude that $\theta$ is continuous by [19, Proposition 6] (whose argument is recalled in the proof of Proposition 8.3), and hence so is $Dw=\gamma(\theta)$. ∎
@@ 8.5.2. The case of a closed curve
Lemma 8.5.2 allows us to use [27, Theorem 6.5], which we recall for the reader’s convenience.
@@
Here and in what follows we implicitly identify a map $v\colon\Omega\to\mathbb{RP}^{1}=\mathbb{S}^{1}/\{\pm 1\}$ with any lifting $\tilde{v}\colon\Omega\to\mathbb{S}^{1}$ such that $v=\{\pm\tilde{v}\}$. In [27, Theorem 6.5], this result is stated for the map $\tilde{v}^{2}$ which uniquely determines $v$ and still satisfies (8.15). Theorem 8.16 is a direct translation of that statement, taking into account the aforementioned implicit identification.
@@ 9. Solutions of degenerate equations
In this final section, we show a stronger and more precise version of Theorem 1.2, Theorem 9.1. First of all, let us introduce the degeneracy set $\mathcal{D}$ in full generality. Following [22, 37], for a strictly monotone $G\in C^{0}(\mathbb{R}^{2},\mathbb{R}^{2})$ we define its degeneracy set $\mathcal{D}\doteq\mathcal{D}_{-}\cap\mathcal{D}_{+}$ which can be interpreted as the set of points where the symmetric parts of $DG$ and $(DG)^{-1}$ both have zero eigenvalues. More precisely, the sets $\mathcal{D}_{\pm}$ are:
The latter is also equivalent to $G(x)\in\mathcal{D}_{-}^{c}(G^{-1})$, justifying its interpretation as local ellipticity from above. We are ready to state our main result of this section:
Theorem 9.1.
Let $G\in C^{0}(\mathbb{R}^{2},\mathbb{R}^{2})$ fulfill (2.1), and assume that each connected component of $\mathcal{D}=\mathcal{D}_{-}\cap\mathcal{D}_{+}$ has image through the graph map
$\mathbb{R}^{2}\ni x\mapsto\left(\begin{array}[]{c}x\\ G(x)\end{array}\right)\in\mathbb{R}^{2\times 2}\,$,
contained in a $C^{1}$ curve. If all but a finite number of these components are
Crucial to our proof is the following result of [36], which shows that blowup limits of solutions to (1.1) are either affine or take values into $\mathcal{D}$. Recalling from Proposition 3.2 the correspondence between (1.1) and (2.3), there is a direct reformulation in terms of differential inclusions, which will be more convenient for us.
@@
Let $K\subset\mathbb{R}^{2\times 2}$ satisfy (2.4), $w\in\operatorname{Lip}(B_{1},\mathbb{R}^{2})$ solve $Dw\in K$ a.e. in $B_{1}$, and $x_{0}\in B_{1}$. Any blowup limit $w_{\infty}\in\mathcal{B}(w)(x_{0})$ is either affine or satisfies $Dw_{\infty}\in\mathcal{K}_{*}$ a.e., where
@@
In this section, we combine this property of blowup limits with Theorems 6.2-8.1, Corollary 7 and structural assumptions on the degenerate set $\mathcal{D}$, or equivalently $\mathcal{K}_{*}$, in order to deduce partial regularity properties of (2.3), or equivalently (1.1), obtaining in particular Theorem 9.1.
@@ 9.1. Regularity threshold
Thanks to Theorem 6.2, it suffices to show that the set $\operatorname{Reg}(w)$ defined in Definition 6 is open. Let $\varepsilon=\varepsilon_{*}(K)/2>0$. By Theorem 4.1, if $x\in\operatorname{Reg}(w)$ then there exists $\delta>0$ such that $\operatorname{diam}([Dw](B_{\delta}(x)))\leq\varepsilon$, so $w$ is $C^{1}$ in $B_{\delta}(x)$ by definition (9.4) of $\varepsilon_{*}(K)$, hence $B_{\delta}(x)\subset\operatorname{Reg}(w)$. ∎
Thanks to Theorem 9.2 and Theorem 4.1, we can relate the regularity threshold $\varepsilon_{*}$ of $K$ with the regularity thresholds of the connected components of $\mathcal{K}_{*}$.
@@
If the infimum is zero, there is nothing to show. Assume therefore that it is positive, and fix $0<\varepsilon<\inf\{\varepsilon_{*}(\mathcal{C})\}$. Let $w\colon B_{1}\to\mathbb{R}^{2}$ satisfy $Dw\in K$ a.e. and $\operatorname{diam}([Dw](B_{1}))\leq\varepsilon$. This pointwise constraint is preserved under blowup: for any $x_{0}\in B_{1}$ and $w_{\infty}\in\mathcal{B}(w)(x_{0})$ we have $\operatorname{diam}([Dw_{\infty}])(B_{1})\leq\varepsilon$. By Theorem 9.2, the blowup limit $w_{\infty}$ is either affine or satisfies $Dw_{\infty}\in\mathcal{K}_{*}$ a.e., hence $Dw_{\infty}\in\mathcal{C}$ a.e. for some connected component $\mathcal{C}$ of $\mathcal{K}_{*}$ by Theorem 2.4. By definition of $\varepsilon_{*}(\mathcal{C})$ this implies $w_{\infty}\in C^{1}(B_{1},\mathbb{R}^{2})$, hence $x_{0}\in\operatorname{Reg}(w)$ by Definition 6. Thus $\operatorname{Reg}(w)=B_{1}$ and $\varepsilon_{*}(K)\geq\varepsilon$. ∎
@@
Under this assumption, every connected component $\mathcal{C}$ of $\mathcal{K}_{*}$ is either a point, or satisfies $\mathcal{C}\subset\Gamma=\gamma(I)$ for some $I=[a,b]$ or $\mathbb{R}/L{\mathbb{Z}}$ and $\gamma\colon I\to\Gamma$ a $C^{1}$ homeomorphism with $|\gamma^{\prime}|>0$. In the latter case, since $\mathcal{C}$ is compact and connected and $\gamma$ is a homeomorphism, we must in fact have $\mathcal{C}=\gamma(J)$ for some compact interval $J\subset I$. So (9.5) is equivalent to every connected component $\mathcal{C}$ of $\mathcal{K}_{*}$ being either a point or a compact $C^{1}$ curve, which implies $\varepsilon_{*}(\mathcal{C})>0$ by Proposition 8.4. In fact, we will see in §9.2 that under assumption (9.5), the property $\varepsilon_{*}(K)>0$ implies a much stronger conclusion than that of Lemma 9.1.
@@
This can only happen if $a=b$, and hence $h$ fulfills (3.10). Since $h$ is one-homogeneous, $w_{\alpha}(z)\doteq\alpha w(z)$ solves (3.8) for $h$ as in (9.6) for all $\alpha\in\mathbb{R}$. Moreover $\operatorname{Sing}(w_{\alpha})=\{0\}$, regardless of how small $\alpha$ is, showing that no $\varepsilon$-regularity theorem can hold for the PDE (3.8) for such $h$. In other words, the regularity threshold (9.4) of the set $K=\{(a,h(a))\colon|a|\leq 1\}$ is $\varepsilon_{*}(K)=0$. The problem in this case is, in fact, fully degenerate: for $h$ as in (9.6), passing to the equivalent formulation in terms of a monotone field $G$ as in (1.1) (see Propositions 3.2-3.2), we find that $\mathcal{D}=\mathcal{D}_{+}\cap\mathcal{D}_{-}=\mathbb{\mathbb{R}}^{2}$, or equivalently $K=\mathcal{K}_{*}$. To see this, one may use the calculations in the proof of [37, Theorem 1.5], which yield $F(\mathbb{S}^{1})\subset\mathcal{D}$, where $F$ is the homeomorphism
$F:\mathbb{C}\to\mathbb{C},\quad F(z)\doteq\frac{h(z)+\bar{z}}{2}.$
Since $h$ is 1-homogeneous, the same calculations show that $F(t\mathbb{S}^{1})\subset\mathcal{D}$ for all $t>0$, and, since $F(\mathbb{C})=\mathbb{C}$ and $\mathcal{D}$ is closed, we get that $\mathcal{D}$ must be, in this case, the whole plane.
9.2. Characterization of blowup limits
Since blowup limits $w_{\infty}$ are entire solutions of $Dw_{\infty}\in\mathcal{C}$ a.e. for some connected component $\mathcal{C}$ of $\mathcal{K}_{*}$, under assumption (9.5) we know by Theorem 8.1 that they are $C^{1}$ away from at most two singular points. This statement is valid for any entire solution, but blowup limits have the additional property that their Hessian determinant vanishes away from the origin by Corollary 7. Using this, we establish that they have in fact at most one singularity.